OpenStudy (anonymous):

Factor completely 25x^2 – 36

OpenStudy (anonymous):

difference of squares: a^2-b^2=(a+b)(a-b) your a=5x your b=6

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

what do you mean, what?

OpenStudy (anonymous):

im just confused...

OpenStudy (anonymous):

your are dealing with a difference of squares, in general a difference of squares has this form: a^2+b^2 and it factors to : (a+b)(a-b) In you problem we have (5x)^2-(6)^2